# Dynamic Rank Adjustment for Accurate and Efficient Neural Network Training

### v2

@@ 4.2 Model Rank Scheduling
First, re-write each matrix as a summation of rank-1 matrices such that $\mathbf{W}_{0}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}$ and $\mathbf{A}\mathbf{B}^{\top}=\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. Consequently, $\mathbf{W}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}+\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. The minimum rank of $\mathbf{W}$ is rkr-k if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ eliminate $k$ rank-1 components of $\mathbf{W}_{0}$. On the other hand, the maximum rank of $\mathbf{W}$ is $r+k$ if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ are orthogonal to all $k$ components of $\mathbf{W}_{0}$. ∎
To connect theory and practice, we compare weight ranks under different rank schedules in Table 2. We perform CIFAR-10 (ResNet-20) training with the learning rate decaying at epoch 100. The rightmost three columns show $\lambda$ for layers 0, 10, and 18 (near the input, middle, and output). In case 1 (conventional low-rank), $\lambda_{0}$ and $\lambda_{18}$ are very large, $\infty$ denotes values $10^{6}$, indicating widely spread spectra and effective rank loss. As full-rank training is applied later, the spectra narrow (smaller $\lambda$ values). These results demonstrate that the full-rank budget should be placed toward the end of the high-noise regime.
Now, let us discuss the impact of full-rank training in the low-noise regime. Once the learning rate decays, the rank of the model weights does not change dramatically due to the reduced magnitude of updates. Instead, what matters more is the extent to which low-rank training pushes the model away from the full-rank convergence point. The smaller the learning rate, the less significant the difference between low-rank and full-rank training becomes.
Proposition 2. As the learning rate decays, the gap between the full-rank model update and the low-rank model update is likely to decrease.
Proof.
The gradient of low-rank model weights, $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ can be written as $\nabla f(\mathbf{A}\mathbf{B}^{\top})\mathbf{B}\in\mathbb{R}^{m\times k}$ and $\nabla f(\mathbf{A}\mathbf{B}^{\top})^{\top}\mathbf{A}\in\mathbb{R}^{n\times k}$, respectively. Therefore, the reconstructed model can be written as follows.
@@
$=\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$
$\hskip 14.22636pt-\eta\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{B}_{t}^{\top}-\eta\mathbf{A}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t})$
$\hskip 14.22636pt+\eta^{2}\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t}).$
@@
Then, the difference between the full-rank update $\nabla f(\mathbf{W}_{t})$ and the low-rank update, $d$ is defined as follows.
$d_{t}:=\left\|\nabla f(\mathbf{W}_{t})-\left(-\mathbf{X}_{t}-\mathbf{Y}_{t}+\eta\mathbf{Z}_{t}\right)\right\|_{F}$ (2)
@@
$d_{t}$ $\leq\left\|\nabla f(\mathbf{W}_{t})\right\|_{F}$
$\hskip 14.22636pt+\left\|\mathbf{X}_{t}\right\|_{F}+\left\|\mathbf{Y}_{t}\right\|_{F}+\eta\left\|\mathbf{Z}_{t}\right\|_{F}.$ (3)
@@
$d_{t}$ $\leq\|\nabla f(\mathbf{W}_{t})\|_{F}\left(1+\|\mathbf{A}_{t}\|_{2}^{2}+\|\mathbf{B}_{t}\|_{2}^{2}\right)$
$\hskip 14.22636pt+\eta\|\nabla f(\mathbf{W}_{t})\|_{F}\left(\|\nabla f(\mathbf{W}_{t})\|_{2}\cdot\|\mathbf{A}_{t}\|_{2}\cdot\|\mathbf{B}_{t}\|_{2}\right).$
According to the bound above, the maximum difference decreases as $\eta$ decays. ∎
Note that, for clarity, we omit $\mathbf{W}_{0}$ from the analysis. Incorporating $\mathbf{W}_{0}$ into each step of the proof to recover LoRA-style low-rank training is straightforward. The above analysis shows that full-rank training should be performed as early as possible in the low-noise regime to minimize $d_{t}$. In general, most popular learning rate schedules in neural network training monotonically decrease the learning rate over time. Therefore, performing full-rank training earlier tends to result in a smaller accumulated difference, $\sum_{t}d_{t}$, during the rest of the training epochs.
@@ 4.3 Unified Rank Adjustment Framework
Based on the two key insights discussed above, we build up a general model rank adjustment framework as shown in Figure 2. The training begins in a low-rank space. The model’s rank is adjusted twice throughout the training, increased in a high-noise regime and then decreased in a low-noise regime. In our empirical study, we observed reasonably good performance when $I$ is set to the middle of high-noise regime and $D$ to the middle of low-noise regime. For example, given $E=150$, if learning rate decays at epoch $100$, we recommend beginning with $I=50$ and $D=125$ and then fine-tune these values to minimize training time while maintaining model accuracy. We will include an ablation study to analyze how the choices of $I$ and $D$ affect model accuracy and computational cost in the following section.
@@ 5 Experiments
Experimental Settings – We evaluate dynamic rank training framework on three computer vision benchmarks: CIFAR-10 [19] (ResNet20), CIFAR-100 (Wide-ResNet28-10), and TinyImageNet [24] (ResNet50), and eight natural language processing (NLP) benchmarks in GLUE [29] (DeBERTaV3-base [9]). All experiments are conducted on a GPU server equipped with two NVIDIA RTX 4090 GPUs. Each experiment is repeated at least twice, and we report the average accuracy along with the standard deviation. Detailed hyperparameter settings are provided in the Appendix.
@@ 5.1 Comparative Study
Computer Vision Benchmarks – To evaluate the performance of our dynamic-rank training strategy, we compare it with several low-rank reparameterization methods: SVD [14], Tucker [18], and CP [20] decompositions. We also compare it with TKD-CPD [25], a recently proposed low-rank training method that jointly utilizes Tucker and CP decompositions. Table 3 shows the results. The Rank Ratio $\rho$ indicates the ratio of the reparameterized weight rank to the original model’s maximum rank. Comp. refers to the average proportion of trainable parameters during training. Across all benchmarks, low-rank training leads to noticeable drops in accuracy. While increasing $\rho$ to 0.75 narrows the gap, it still falls short of full-rank performance. In contrast, dynamic-rank training matches full-rank accuracy while maintaining a similar computational cost to the $\rho=0.75$ setting. While TKD-CPD dramatically reduces the computational cost, it yields the lowest model accuracy among all the low-rank methods. However, when dynamic-rank training is applied to TKD-CPD, its accuracy is restored to a level comparable to full-rank training. These results suggest that the dynamic-rank training framework is readily applicable to a variety of modern machine learning applications.
Method Comp. CoLA MNLI MRPC QNLI QQP RTE SST-2 SST-B Average
Full Fine-Tuning $1.000$ $0.6750$ $0.9018$ $0.8995$ $0.9397$ $0.9222$ $0.8375$ $0.9587$ $0.9097$ $0.8764$
LoRA $0.014$ $0.6333$ $0.8874$ $0.8799$ $0.9192$ $0.8922$ $0.7978$ $0.9461$ $0.9021$ $0.8618$
AdaLoRA $0.017$ $0.6599$ $0.8891$ $0.8431$ $0.9183$ $0.8891$ $0.8086$ $0.9477$ $0.8943$ $0.8564$
Dynamic Rank (SVD) $0.115$ $\mathbf{0.6752}$ $\mathbf{0.8959}$ $\mathbf{0.8886}$ $\mathbf{0.9264}$ $\mathbf{0.9062}$ $\mathbf{0.8267}$ $\mathbf{0.9481}$ $\mathbf{0.9087}$ $\mathbf{0.8714}$
Table 4: Test accuracy comparison with NLP model (DeBERTaV3-base) and GLUE benchmark.
@@
Natural Language Processing Benchmarks – We validate dynamic-rank training framework by fine-tuning the DeBERTaV3-base on 8 datasets from GLUE benchmark [29]. In fine-tuning, there is no high-noise regime, as the learning rate is typically set to a small value and remains in the low-noise regime throughout training. Therefore, in these experiments, we place full-rank epochs at the beginning of fine-tuning.
Table 4 reports the results. We compare our method with LoRA [10], the most widely used low-rank fine-tuning method, and AdaLoRA [33], a representative variant. All low-rank methods greatly reduce compute versus full fine-tuning. LoRA is the most efficient but shows a noticeable accuracy drop. AdaLoRA mitigates the drop but still does not match the full fine-tuning accuracy. By contrast, dynamic-rank training achieves accuracy comparable to that of full fine-tuning. Although the computational cost is roughly $7$ times AdaLoRA’s cost but still far cheaper than full fine-tuning, making it a practical choice for NLP applications.
Dynamic-rank Training with Regularization – Our dynamic-rank training strategy is readily applicable to regularization methods designed to restore the reduced effective rank.
Table 5 shows CIFAR-10 benchmark results with three regularization methods, soft orthogonality (SO) [32], double soft orthogonality (DSO) [1], and SRIP+ [17]. These regularizers effectively improve the accuracy of SVD-based low-rank training. When combined with dynamic-rank training, they yield consistent accuracy gains across all regularizers. Therefore, the two are complementary, and their combination achieves high accuracy at low computational cost. However, these regularizers consume a large amount of memory space to compute Gram matrices, and the overall memory footprint may substantially increase as they are combined with dynamic-rank training.
Parameter Heatmap Comparison – Figure 4 shows a parameter value comparison between low-rank training and dynamic-rank training. We draw heatmaps with 4-D weight tensors of size $3\times 3\times 64\times 64$ obtained from the largest convolution layer in ResNet20. The tensors are reshaped to $192\times 192$ 2-D matrices. To show the difference more clearly, we reconstructed the tensor from the low-rank reparameterized weights using the 50 singular vectors corresponding to the smallest singular values. The low-rank model’s map exhibits blocky artifacts, whereas the dynamic-rank model’s map does not. The presence of blocky artifacts implies the existence of a null space in certain directions of the feature space, meaning that variations in the input along those directions do not affect the output. This result is well aligned with the model accuracy comparisons shown in other tables. Consequently, we conclude that the dynamic-rank training framework well restores the effective rank of the weight matrix, thereby improving model accuracy.
@@
5.2 Ablation Study
In our empirical study, we found that dynamic-rank training performs well when $D$ and $I$ are set to allocate a similar number of full-rank epochs to both the high-noise and low-noise regimes. We now conduct a simple ablation study to examine how the number of full-rank epochs affects model accuracy. Table 6 presents the results. SVD-based low-rank training is evaluated on the CIFAR-10 and CIFAR-100 benchmarks. For convenience, we define $\phi=(D-I)/E$ as the ratio of full-rank epochs to the total training budget. When $\phi=0.0$, it corresponds to conventional low-rank training; when $\phi=1.0$, it becomes full-rank training. On both benchmarks, accuracy starts to drop when $\phi$ goes below 0.5, particularly in the range $\phi\in[0.3,0.5]$. We therefore recommend starting with $\phi=0.5$ and adjusting downward as needed.
@@ 5.2 Ablation Study
Setting Rank Comp. CIFAR-10 CIFAR-100
Ratio $\rho$
$\phi=0.1$ 0.25 0.36 $90.54\pm 0.3\%$ $76.29\pm 0.1\%$
$\phi=0.3$ 0.51 $91.43\pm 0.1\%$ $77.03\pm 0.4\%$
$\phi=0.5$ 0.64 $92.01\pm 0.3\%$ $78.61\pm 0.2\%$
$\phi=0.7$ 0.79 $92.08\pm 0.4\%$ $78.75\pm 0.1\%$
$\phi=0.9$ 0.93 $92.14\pm 0.2\%$ $78.63\pm 0.3\%$
$\phi=0.1$ 0.5 0.61 $91.28\pm 0.1\%$ $76.84\pm 0.1\%$
$\phi=0.3$ 0.69 $91.57\pm 0.1\%$ $77.51\pm 0.2\%$
$\phi=0.5$ 0.78 $92.11\pm 0.2\%$ $78.41\pm 0.3\%$
$\phi=0.7$ 0.87 $92.15\pm 0.2\%$ $78.39\pm 0.1\%$
$\phi=0.9$ 0.96 $92.11\pm 0.1\%$ $78.47\pm 0.3\%$
Table 6: Ablation study on how $\phi=(D-I)/E$ affects the performance of dynamic-rank training. Low-Rank SVD is used for all experiments. When $\phi=0.0$, it becomes the conventional low-rank training. In contrast, when $\phi=1.0$, it becomes full-rank training. The $I$ and $D$ are set to perform the same number of full-rank epochs in the high-noise and low-noise regimes.

### v3

@@ 4.2 Model Rank Scheduling
The gradient of adaptor weights, $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are written as $\nabla f(\mathbf{A}\mathbf{B}^{\top})\mathbf{B}\in\mathbb{R}^{m\times k}$ and $\nabla f(\mathbf{A}\mathbf{B}^{\top})^{\top}\mathbf{A}\in\mathbb{R}^{n\times k}$, respectively. So, the reconstructed model is:
@@
$=\mathbf{A}_{t}\mathbf{B}_{t}^{\top}-\eta\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{B}_{t}^{\top}-\eta\mathbf{A}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t})+\eta^{2}\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t}).$
@@
Then, the difference between the full-rank update $\nabla f(\mathbf{W}_{t})$ and the low-rank update, $d$ becomes:
$d_{t}:=\left\|\nabla f(\mathbf{W}_{t})-\eta\left(-\mathbf{X}_{t}-\mathbf{Y}_{t}+\eta\mathbf{Z}_{t}\right)\right\|_{F}$ (2)
@@
$d_{t}$ $\leq\left\|\nabla f(\mathbf{W}_{t})\right\|_{F}+\eta\left\|\mathbf{X}_{t}\right\|_{F}+\eta\left\|\mathbf{Y}_{t}\right\|_{F}+\eta^{2}\left\|\mathbf{Z}_{t}\right\|_{F}.$ (3)
@@
$d_{t}$ $\leq\|\nabla f(\mathbf{W}_{t})\|_{F}\left(1+\eta\|\mathbf{A}_{t}\|_{2}^{2}+\eta\|\mathbf{B}_{t}\|_{2}^{2}\right)$
$\hskip 14.22636pt+\eta^{2}\|\nabla f(\mathbf{W}_{t})\|_{F}\left(\|\nabla f(\mathbf{W}_{t})\|_{2}\cdot\|\mathbf{A}_{t}\|_{2}\cdot\|\mathbf{B}_{t}\|_{2}\right).$
According to the bound above, the difference $d_{t}$ is expected to decreases as $\eta$ decays. ∎
For clarity, we omit $\mathbf{W}_{0}$ from the analysis. Incorporating $\mathbf{W}_{0}$ into each step of the proof to recover LoRA-style low-rank training is straightforward. The above analysis shows that full-rank training should be performed as early as possible in the low-noise regime to minimize $d_{t}$. In general, most popular learning rate schedules monotonically reduce the learning rate over time. Therefore, scheduling full-rank epochs earlier tends to result in a smaller accumulated difference, $\sum_{t}d_{t}$, over the remaining training epochs.
@@ 4.3 Unified Rank Adjustment Framework
Based on the two key insights discussed above, we build up a general model rank adjustment framework as shown in Figure 2. The training begins in a low-rank space. The model’s rank is adjusted twice throughout the training, increased in a high-noise regime and then decreased in a low-noise regime. In our empirical study, we observed reasonably good performance when $I$ is set to the middle of high-noise regime and $D$ to the middle of low-noise regime. For example, given $E=150$, if learning rate decays at epoch $100$, we recommend beginning with $I=50$ and $D=125$ and then fine-tune these values to minimize training time while maintaining model accuracy. The pseudocode is provided in Appendix 7.1.
@@ 5 Experiments
Experimental Settings – We evaluate dynamic-rank training framework on three computer vision benchmarks: CIFAR-10 (krizhevsky2009learning, ) (ResNet20), CIFAR-100 (Wide-ResNet28-10), and TinyImageNet (tiny-imagenet, ) (ResNet50), and natural language processing (NLP) datasets in GLUE (wang2018glue, ) (DeBERTaV3-base (he2021debertav3, )) benchmark. All experiments were conducted on a GPU server with two NVIDIA RTX 4090 GPUs. Each experiment was repeated at least twice, and we report the mean accuracy with standard deviation. The detailed hyperparameter settings are presented in the Appendix.
@@ 5.1 Comparative Study
Computer Vision Benchmarks – To evaluate the performance of our dynamic-rank training strategy, we compare it with several low-rank reparameterization methods, including SVD, Tucker, and CP decompositions. We also compare it with TKD-CPD (phan2020stable, ), a recently proposed low-rank training method that jointly utilizes Tucker and CP decompositions. Table 4 shows the results. See Appendix for $I$ and $D$ settings. The Rank Ratio $\rho$ indicates the ratio of the reparameterized weight rank to the original weight rank. Comp. refers to the average proportion of trainable parameters during training. Across all benchmarks, low-rank training leads to noticeable accuracy drops. Increasing $\rho$ to 0.75 narrows the gap but still falls short of full-rank performance. In contrast, dynamic-rank training achieves full-rank accuracy at a computational cost similar to the $\rho=0.75$ setting. TKD-CPD dramatically reduces computation but yields the lowest accuracy among low-rank methods. Applying dynamic-rank training to TKD-CPD restores its accuracy to a level comparable to full-rank training.
Table 4: Comparison of CV benchmarks across various low-rank training methods. The dynamic-rank method consistently improves accuracy while effectively reducing computational cost.
Acc. Comp. Acc. Comp. Acc. Comp.
SVD $0.50$ $91.19\pm 0.1\%$ $0.57$ $73.57\pm 0.1\%$ $0.56$ $54.90\pm 0.1\%$ $0.79$
$0.75$ $91.56\pm 0.1\%$ $0.84$ $76.90\pm 0.1\%$ $0.84$ $56.16\pm 0.1\%$ $0.92$
$0.75$ $91.50\pm 0.1\%$ $0.79$ $79.68\pm 0.2\%$ $0.77$ $61.07\pm 0.7\%$ $0.87$
CP $0.50$ $88.31\pm 0.1\%$ $0.31$ $65.55\pm 0.1\%$ $0.30$ $42.05\pm 0.1\%$ $0.65$
$0.75$ $88.90\pm 0.1\%$ $0.59$ $65.32\pm 0.1\%$ $0.57$ $43.12\pm 0.1\%$ $0.78$
0.75 $89.62\pm 0.1\%$ 0.35 $60.22\pm 0.1\%$ $0.31$ $34.84\pm 0.1\%$ 0.66
Table 5: Comparison of NLP benchmark performance on GLUE (DeBERTaV3-base). The dynamic-rank method achieves the best accuracy while maintaining substantially lower computational cost.
Method CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B Avg Comp.
Full Fine-Tuning $0.675$ $0.9018$ $0.8995$ $0.9397$ $0.9222$ $0.8375$ $0.9587$ $0.9097$ $\% 1.0001.0$$0$
LoRA (hu2022lora, ) $0.6333$ $0.8874$ $0.8799$ $0.9192$ $0.8922$ $0.7978$ $0.9461$ $0.9021$ $97.2\%$ $\mathbf{0.014}$
AdaLoRA (zhang2023adalora, ) $0.6529$ $0.8891$ $0.8431$ $0.9183$ $0.8891$ $0.8086$ $0.9495$ $0.8943$ $97.1\%$ $0.017$
DoRA (liu2024doraweightdecomposedlowrankadaptation, ) $0.6579$ $0.8929$ $0.8774$ $0.9176$ $0.8924$ $0.7617$ $0.9506$ $0.8970$ $97.1\%$ $0.014$
SLTrain (han2024sltrainsparsepluslowrank, ) $0.6652$ $0.8913$ $0.8848$ $0.9174$ $0.8934$ $0.7833$ $0.9461$ $0.8987$ $97.6\%$ $0.018$
LoRA-GA (wang2024loragalowrankadaptationgradient, ) $0.6409$ $0.8899$ $0.8602$ $0.9156$ $0.8904$ $0.7617$ $0.9472$ $0.8962$ $96.4\%$ $0.014$
SparseLoRA (khaki2025sparseloraacceleratingllmfinetuning, ) $0.6675$ $0.8898$ $0.8898$ $0.9148$ $0.8900$ $0.7653$ $0.9438$ $0.8960$ $97.3\%$ $0.014$
Dynamic Rank (SVD) $\mathbf{0.6752}$ $\mathbf{0.8959}$ $\mathbf{0.8886}$ $\mathbf{0.9264}$ $\mathbf{0.9062}$ $\mathbf{0.8267}$ $\mathbf{0.9472}$ $\mathbf{0.9087}$ $\mathbf{99.0\%}$ $0.115$
@@
Natural Language Processing Benchmarks – We validate dynamic-rank training framework by fine-tuning the DeBERTaV3-base on 8 datasets from GLUE benchmark (wang2018glue, ). In fine-tuning, there is no high-noise regime, as the learning rate is set to a small value and remains in the low-noise regime throughout training. Therefore, in these experiments, we place full-rank epochs at the beginning of fine-tuning.
We compare our dynamic-rank method with the following SOTA low-rank fine-tuning methods: LoRA, AdaLoRA, DoRA, SLTrain, LoRA-GA, and SparseLoRA. AdaLoRA follows a similar principle to our dynamic-rank method, adjusting the rank at run time. However, it requires costly SVD operations during training, whereas our method avoids explicit decomposition of the weight matrices. DoRA and LoRA-GA have exactly the same computational cost as LoRA because they focus on the initialization of adaptor weights. SparseLoRA has fewer frozen parameters than LoRA, but the number of trainable parameters is the same.
Table 5 reports the fine-tuning performance on GLUE tasks. The Avg column shows the relative performance with respect to full fine-tuning, averaged over the 8 tasks. All the SOTA methods substantially reduce computational cost compared to full fine-tuning. LoRA is the most efficient but suffers from a noticeable drop in accuracy. Other methods alleviate this drop, however the gap from full fine-tuning accuracy remains non-negligible. By contrast, the dynamic-rank training achieves accuracy comparable to the full fine-tuning. This comparison demonstrates that, although the dynamic-rank scheme slightly increases the computational cost compared to LoRA, it effectively restores the effective rank of the model weights, thereby improving the fine-tuning accuracy.
@@
Dynamic-rank Training with Regularization – Our dynamic-rank training strategy is readily applicable to regularization methods designed to restore the reduced effective rank. Table 6 shows CIFAR-10 benchmark results with three regularization methods, soft orthogonality (SO) (xie2017all, ), double soft orthogonality (DSO) (bansal2018can, ), and SRIP+ (kim2022revisiting, ). These regularizers effectively improve the accuracy of SVD‑based low‑rank training. When combined with dynamic‑rank training, they yield consistent accuracy gains across all regularizers. Therefore, the two are complementary, and their combination achieves high accuracy at low computational cost. However, these regularizers consume a large amount of memory space to compute Gram matrices, and the overall memory footprint may substantially increase as they are combined with dynamic-rank training.
@@
Parameter Heatmap Comparison – Figure 4 shows a parameter value comparison between low-rank and dynamic-rank trainings. We draw heatmaps with 4-D weight tensors of size $3\times 3\times 64\times 64$ obtained from the largest convolution layer in ResNet20. The tensors are reshaped to $192\times 192$ 2-D matrices. To show the difference more clearly, we reconstructed the tensor from the low-rank reparameterized weights using 50 singular vectors corresponding to the smallest singular values. The low-rank map exhibits blocky artifacts, whereas the dynamic-rank map does not. The presence of the artifacts implies the existence of a null space in certain directions of the feature space, meaning that variations in the input along those directions do not affect the output. This result is well aligned with the model accuracy comparisons. Therefore, we conclude that the dynamic-rank method well restores the effective rank of the model, thereby better exploiting representation capacity.